OpenStudy (anonymous):

find (g*f)(2) for the functions below f(x)=x+1, g(x)=1/x^2

OpenStudy (anonymous):

(g*f)(2) = g(2) * f(2) = 2+1÷ 1/4 = 3÷ 1/4 = 3*4 = 12